OpenStudy (anonymous):

Baker corporation employed 3670 people. If this number, 1670 had bachelor degrees, 899 had master degrees, and the remainder had no college degrees. What percentage of the employees had no college degree?

OpenStudy (anonymous):

3670 - (1670 + 899) = 3670 - 2569 = 1101 <-- number of people with no college degrees 1101/3670 = 0.3 = 30% <== 30% had no degree